# CAFEs: Cable-driven Collaborative Floating End-Effectors for Agriculture Applications

### v1

@@ I INTRODUCTION
Robotic systems in agriculture primarily rely on Unmanned Ground Vehicles (UGVs) for harvesting [1, 2, 3] and Unmanned Aerial Vehicles (UAVs) for monitoring and yield estimation [4, 5, 6, 7, 8]. While these systems effectively navigate complex farm environments, they depend on intricate localization methods [9], struggle with scalability, and require significant coordination for large-scale deployment.
Cable-driven robots offer an alternative approach by providing precise positioning without complex navigation [10, 11, 12]. Systems like SkyCam [13] and CU-brick [14] showcase their advantages in structured environments. However, traditional cable-driven robots are constrained by their limited number of end-effectors, making multi-robot coordination challenging [15]. Although reconfigurable cable robot [16, 17, 18] improve workspace flexibility, they introduce hardware and control complexities due to multiple interacting cables.
Suspended cable-driven systems can avoid some of these limitations, but they often suffer from restricted workspaces near the ground [19] and require complex real-time control to mitigate oscillations [20, 21, 22]. Additionally, when extending these systems to tens or hundreds of meters, challenges such as cable flexing, uneven tension, and reduced control stability emerge. Flexing under variable loads can introduce positioning errors in the centimeter range, while differential sagging affects precision, limiting their use in large-scale farming.
@@
The challenges of using mobile and cable-driven systems highlight the need for a solution that can cover large areas and allow multiple robots to work together efficiently. We propose a cable-driven system for collaborative end effectors (CAFEs) that adresses these problems by introducing a cable-driven setup that allows several robotic arms to work together without requiring complex navigation or workspace rearrangement as shown in Figure 1. The proposed system leverage the advantages of cable driven robots whilst still enabling cooperation between multiple end effectors (e.g. for tasks such as harvesting). We propose a cable driven system where there are passive cables but also a commonly shared roller-driven cable with runs the length of the floating system. By providing each floating platform with the ability to independently clamp on and off (in either direction of the cable) we maintain the potential for independent operation whilst enabling far simpler synchronzied control as they can be moved by the same rollder-driven cable. Because the design allow multiple robots traversing on the the set of cable, such terrain-independent but still coordinated advantage is able to address the problem in both UGVs and UAVs. Also, the robotic arms can travel in the air with low energy consumption due to the shared weight on the cables. Furthermore, asthe actuation is provided by the shared moving cable each platform remains light weight and requires only clamping control. Specifically the contributions of this work include:
@@ V CONCLUSIONS
The development and testing of the cable-driven Collaborative Agricultural Floating End-effectors (CAFEs) system demonstrate its potential for large-scale agricultural applications. By integrating collaborative capabilities, a fast-clamping mechanism, and dynamic scalability, the system overcomes limitations of traditional cable-driven robots restricted to a single end-effector.
Experiments confirmed high positioning accuracy, stable cooperative motion, and repeatability, while the validated dynamic model for tension distribution and cable sag compensation ensures stability in larger deployments. The results highlight the feasibility of CAFEs for harvesting, monitoring, and object transport, improving operational efficiency.
Future work will focus on enhancing control algorithms, optimizing hardware, and refining motion planning to address multi-robot coordination and long-term deployment challenges. Further development of anti-oscillation damping using the dynamic model will improve system stability, positioning CAFEs as a key technology in agricultural automation.

### v2

@@ I INTRODUCTION
Robotic systems in agriculture primarily rely on Unmanned Ground Vehicles (UGVs) for harvesting [1, 2, 3] and Unmanned Aerial Vehicles (UAVs) for monitoring and yield estimation [4, 5, 6, 7, 8]. While these systems effectively navigate complex farm environments, they rely on GPS, Real-Time Kinematic (RTK) positioning, Simultaneous Localization and Mapping (SLAM), or visual-inertial odometry for self-localization [9]. However, these localization methods can be intricate, adding to the challenges of cost, scalability, and coordination for large-scale deployment. A well-designed cable-driven mechanism could offer a viable alternative, reducing localization dependency while maintaining efficient coverage of agricultural fields.
Cable-driven robots offer an alternative approach by providing precise positioning without complex navigation [10, 11, 12]. Examples like SkyCam [13] and CU-brick [14] showcase their advantages in large spaces without the need for complex navigation, making them useful for agricultural work. However, traditional cable-driven robots are constrained by their scalability due to the limited number of end-effectors.The multi-robot coordination is challenging regarding to the cable interference [15]. Although reconfigurable cable robot [16, 17, 18] improve workspace flexibility, they introduce hardware and control complexities due to multiple interacting cables.
Suspended cable-driven systems can avoid some of these limitations, but they often suffer from limited workspaces near the ground [19] and require complex real-time control to mitigate oscillations [20, 21, 22]. Additionally, when scaling up these systems to tens or hundreds of meters, cable elongation and flexing under variable loads can introduce positioning errors in the centimeter range, while differential sagging affects precision, limiting their use in large-scale farming.
@@
The challenges of using mobile and cable-driven systems highlight the need for a solution that can cover large areas and allow multiple robots to work together efficiently. We propose a cable-driven system for collaborative end effectors (CAFEs) that adresses these problems by introducing a cable-driven setup that allows several robotic arms to work together without requiring complex navigation or workspace rearrangement as shown in Figure 1. The proposed system leverages the advantages of cable-driven robots while enabling cooperation between multiple end-effectors, such as for harvesting tasks. It can be deployed in both greenhouses and outdoor fields by strategically placing multiple anchoring points. The system’s scalability is achieved by adding more CAFEs, incorporating additional cable sets, or extending the cable lengths.
The proposed cable-driven system features both passive cables and a commonly shared roller-driven cable that runs along the length of the floating structure. By providing each floating platform with the ability to independently clamp on and off (in either direction of the cable) we maintain the potential for independent operation whilst enabling far simpler synchronzied control as they can be moved by the same rollder-driven cable. Because the design allow multiple robots traversing on the the set of cable, such terrain-independent but still coordinated advantage is able to address the problem in both UGVs and UAVs. Also, the robotic arms can travel in the air with low energy consumption due to the shared weight on the cables. Furthermore, asthe actuation is provided by the shared moving cable each platform remains light weight and requires only clamping control. Specifically the contributions of this work include:
@@ V CONCLUSIONS
The development and testing of the cable-driven Collaborative Agricultural Floating End-effectors (CAFEs) system demonstrate its potential for large-scale agricultural applications by integrating collaborative capabilities, a fast-clamping mechanism, and dynamic scalability, overcoming the limitations of traditional cable-driven robots restricted to a single end-effector. Experiments confirmed high positioning accuracy, stable cooperative motion, and repeatability, while the validated dynamic model for tension distribution and cable sag compensation ensures stability in larger deployments. The results highlight the feasibility of CAFEs for harvesting, monitoring, and object transport, improving operational efficiency. Future work will focus on enhancing control algorithms, optimizing hardware, and refining motion planning to address multi-robot coordination, while further development of anti-oscillation damping will improve system stability, positioning CAFEs as a key technology in agricultural automation.